Use the Intermediate Value Theorem to show that the graph of the function has a zero in the given interval. Round the zero to two decimal places.
f(x) = 2x3 + x2 – x – 5; [1, 2] Please help

Answer by venugopalramana(3286)

Use the Intermediate Value Theorem to show that the graph of the function has a zero in the given interval. Round the zero to two decimal places.
f(x) = 2x3 + x2 – x – 5; [1, 2] Please help
F(1)=2+1-1-5=-3
F(2)=2*2^3+2^2-2-5=16+4-2-5=13
F(X) IS CONTINUOUS IN THIS RANGE AND CHANGES ITS SIGN FROM -VE TO +VE.HENCE IT HAS ZERO IN THIS RANGE
NARROWING THE RANGE TO GET AT THE ZERO..TRY X=(1+2)/2=1.5
F(1.5)=2.5..
THEN...TRY...(1+1.5)/2=1.25 ..F(1.25)=-0.78
THEN ..TRY....(1.25+1.5)/2=1.375....F(1.375)=0.714.......ETC..
WE FIND BY TRIAL THAT AT X=1.32 F(X)=0